Grade 12Mechanics

Consider a system of two identical particles. One of the particles is at rest and the other has an acceleration f. The centre of mass has an acceleration 1) 0. 2) f. 3)f/2. 4) 2f. I also want to know when`s the does the centre of mass has an acceleration 0, f, f/2, 2f

Sakshi
9 Years agoGrade 12

2 Answers

Raman Mishra
ApprovedApproved Tutor Answer9 Years ago
Let acm be the acceleration of the centre of mass of the two particles. Then,
                a_{cm}=\frac{m_{1}a_{1}+m_{2}a_{2}}{m_{1}+m_{2}}
Here since the particles are identical i.e, m1=m2=m(say)
          a_{cm}=\frac{m(a_{1}+a_{2})}{2m}=\frac{a_{1}+a_{2}}{2}               (1)
Taking a1=0(particle at rest) and a2= f, we get
        acm=f/2   Ans.
Also, putting the desired values of acm in eqn(1), we have:
acm =0 when a1+a= 0 ;
acm = f when a1+a= 2f;
acm = f/2 when a1+a= f and
acm = 2f when a1+a= 4f.
Hope you understood dear Sakshi.
Abcd
6 Years ago
[5kg]-------------[2kg]
V. Com=m1v1 + m2v2 /m1+m2
 
It is given that mass of 2 kg is at rest....
.: V. Com=5×7/7
         =5m/sec ...😊